RAMAN SPECTRA OF METHANOL AND ITS SOLUTIONS

\underline{F. H. TUKHVATULLIN}, U. N. TASHKENBAEV and A. JUMABOEV, Department of Physics, Samarkand State University, Samarkand 703004 Uzbekistan.

Raman spectra of C-O vibration for methanol were studied. The band is asymmetric in low-frequency side and besides, the depolarization coefficient within of band is non-monotonously changed. In solutions with CCl_{4} at small concentration of alcohol (less 0,1 mol. fraction) two lines correspond for this vibration. In alcohol-acetonitrile solutions over the whole region of concentration, the splitting of C-O vibration band into components does not take place. With dilution of alcohol the asymmetry of band is decreased. The results were attributed to formation and destruction of intermolecular H - associations.